How much water should be added to 500 l of a 14% syrup solution to get a 10% syrup solution?
.14x + X + .10x = 500
1.24x = 500
x=403.22

Answer by stanbon(75887)

How much water should be added to 500 l of a 14% syrup solution to get a 10% syrup solution?
----
Equation:
water + water = water
0.86*500 + x = 0.90(500+x)
---
Multiply thru by 100 to get:
86*500 + 100x = 90*500+90x
10x = 4*500
x = 200 L (amt of water that must be added)
==================================================
You could also just solve the following:
sugar + sugar = sugar
0.14*500 + 0*x = 0.10(500+x)
Multiply by 100 to get:
14*500 = 10*500+10x
10x = 4*500
x = 200 L (amt of water that must be added)
==================================================
Cheers,
Stan H.
=============